OpenStudy (anonymous):

(e^x)+(3e^y)=3 and (e^2x)-(9e^2y)=6 find the values of x and Y. some please help me out!!

OpenStudy (anonymous):

If e^x=u and e^y=v. Then u+3v=3. and u^2-9v^2=6. Using a^2-b^2=(a+b)(a-b) (u+3v)(u-3v)=6 u+3v=3. so u-3v=2. and u+3v=3.solve these two.

OpenStudy (anonymous):

thank you very much :D

OpenStudy (mertsj):

\[e ^{2x}-9e ^{2y}=(e ^{x}-3e ^{y})(e ^{x}+3e ^{y})=6\]

OpenStudy (mertsj):

But e^x+3e^y=3 so substitute.

OpenStudy (mertsj):

\[(e ^{x}-3e ^{y})(3)=6\]

OpenStudy (mertsj):

\[e ^x-3e ^{y}=2\]

OpenStudy (mertsj):

But from the first equation \[e^x+3e^y=3\]

OpenStudy (mertsj):

Add those two equations.

OpenStudy (mertsj):

\[2e^x=5, e^x=\frac{5}{2}\]

OpenStudy (mertsj):

\[x=\ln (\frac{5}{2})\]

OpenStudy (mertsj):

\[ y = \ln\frac{1}{6}\]

OpenStudy (mertsj):

Forgot to say that.